Find the Laplace transform F(s) L(f(t)) given f(t) = 5e-4 sin(5t) + 2e cos(6t). F(8) =
Answer #1

Here basically we use the shifting property and linearity property of the Laplace transform.

Guven function f(t)= 5ē at Run (5t) + 2ězt comC6t). Want to find the Laplace transform of f(t). le F(s)= L (160). Now the shiThus Lleut regt}= G, (S- (-41) = G, (S+4) 5+ (S+4) S Now let 92 (1) = (ol 67. then Laplace trane formation of 92(1); L {92(1)2824 Cox (61). Since f(t)=584t ein est + 2e then using the linearity of Laplace transformation F(S) = l{}(1} = 5.44eal Rai 5t

Thus we are done!